A bus travel as 60km/hr.how long those it take cover a distance of 500km
1
Expert's answer
2020-12-21T04:02:52-0500

d = 500 km

v = 60 km/hr

"v = \frac{d}{t} \\

t = \frac{d}{v} \\

t = \frac{500}{60} = 8.33 \;hr"

t = 8 hr 20 min